# DisastIR: A Comprehensive Information Retrieval Benchmark for Disaster Management

### v1

@@ Acknowledgments
Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing
@@ Appendix A Structural PDF File Processing Pipeline
All disaster management-related data (in PDF format) is obtained from publicly available sources with no personally identifiable information. Hence, explicit consent was not required. PDF files are collected using googlesearch-python (v1.3.0) and processed with PyMuPDF (v1.24.10) for content extraction. The extracted PDFs are then processed into text chunks through the following steps:
@@
Appendix B Prompt Templates for Query Generation
Prompts for query generation based on disaster management-related passage under different search intents for QA, QAdoc, Twitter, FC, NLI, STS are in Tables 4, 5, 6, 7, 8, and 9.
Appendix C Prompt Templates for Relevance Labeling
This section presents the prompt templates used for LLM-based relevance judgments across six search intents, employing three prompting strategies: Zero-shot Direct Scoring, Chain-of-Thought Decomposed Reasoning, and Multi-Dimensional Attribute Scoring. For QA, QAdoc, and Twitter tasks, we adapt templates from Thomas et al. (2024); Farzi and Dietz (2024), as shown in Table 10. Based on these templates, we design relevance prompts for FC, NLI, and STS tasks, shown in Tables 11, 12, and 13, respectively. For STS, we adopt only Zero-shot Direct Scoring, as our preliminary experiments show it yields higher agreement with human labels (Cohen’s kappa). The estimated cost of generating 9,600 user queries and labeling over 1.3 million query-passage using GPT-4o-mini API is about \$1,400.
Appendix D Additional Analyses of Labeled query-passage pairs
@@
Appendix E LVHL Dataset Construction
@@
Appendix F LVHQ Dataset Construction
We sample 48 domain passages developed in Section 3.3, ensuring one passage per retrieval task and keeping all sampled passages different from those used in developing DisastIR. For each passage, a domain expert in the disaster management field is asked to read the passage and write a realistic user query that reflects a practical information need based on the content, resulting in 48 human-authored queries. The Human expert is given the same instructions for the query written (shown in Tables 4, 5, 6, 7, 8, 9) as those given to LLM to ensure fair comparison. In parallel, for the same set of passages, we also generate 48 queries using LLM in the same way as described in Section 3.4.
@@ Appendix F LVHQ Dataset Construction
Appendix G Information of Evaluated Models and Model Implementation
@@
Appendix H Performance of Evaluated Models
Performance of all evaluated models in all 48 search tasks in DisastIR is shown in Tables 19, 20, 21, 22, and 23.
Appendix I Additional Analyses of Model Performance across 48 Tasks
NV-Embed-v2 achieves the best performance on all QAdoc and NLI-related tasks (See Table 3 and Figure 5 in the main content). However, as shown in Figure 7, its poor results on Twitter-related tasks significantly lower its overall performance in DisastIR. This reflects its limitation in handling the informal, noisy, and contextually ambiguous nature of social media content. Given the importance of Twitter as a real-time, crowd-sourced information source during disasters (Alam et al., 2021; Yin et al., 2024; Lei et al., 2025), this weakness raises concerns about its reliability in real-world disaster response scenarios.
All three models perform poorly on NLI-related tasks, with the best achieving only an average score of 58.73 (Figure 7). Further analysis of outliers in the box plot (See Figure 3 in the main content) reveals that tasks causing significant performance drops consistently involve NLI search intents (Figure 8). This reveals a key limitation of current open-source SOTA retrievers, that they struggle with the complex reasoning required for NLI tasks in disaster contexts. Such limitations may lead to incorrect results or failure to retrieve critical information, which can negatively impact decision-making in disaster situations.
@@ Appendix I Additional Analyses of Model Performance across 48 Tasks
System prompt: You are a search quality rater evaluating overall relevance. Given a query, passage, and sub-scores, provide a final score (0–3): 3 = Perfectly relevant, 2 = Highly relevant, 1 = Related, 0 = Irrelevant. Your output MUST be JSON: {"final_relevance_score": <0-3>}
User prompt: Please rate how relevant the passage is based on the given sub-scores.Query: {query}, Passage: {passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score.
@@
[1] Zero-shot Direct Scoring
@@
User prompt: Given a claim and a passage, score relevance (003333): 00 = Irrelevant, $1$ = Related (no help), $2$ = Relevant (unclear/mixed), $3$ = Direct support/refutation.
@@
[2] Chain-of-Thought Decomposed Reasoning
@@
PHASE 2 – Fine-grained Criterion Scoring:
System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim. Use the following scoring scale (003333): 00 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion.2222 = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
@@
System prompt when Phase 1 = "Yes": You are a rater evaluating evidence for fact-checking. Score ($2$ or $3$): $2$ = Highly relevant (some support/refutation), $3$ = Perfectly relevant (direct support/refutation). Your output MUST be JSON: {"relevance_score": <2_or_3>} System prompt when Phase 1 = "No": You are a rater evaluating evidence for fact-checking. Score (00 or $1$): 00 = Irrelevant, $1$ = Related (no support/refutation). Your output MUST be JSON: {"relevance_score": <0_or_1>}
@@
Passage irrelevant for evidence. Rate its relevance (00 or $1$). Claim: {Query}, Passage: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
[3] Multi-Dimensional Attribute Scoring
PHASE 1 – Sub-criterion Scoring:
System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim according to the following criteria. Use a score (003333): 00 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion. $2$ = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
@@
System prompt: You are a search quality rater evaluating evidence relevance for fact-checking. Given a claim, passage, and sub-scores, provide a final score (003333): $3$ = Perfectly relevant (direct support/refutation), $2$ = Highly relevant (helps verification), $1$ = Related (topic match, no verification aid), 00 = Irrelevant. Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
@@
[1] Zero-shot Direct Scoring
@@
User prompt: Given a premise and hypothesis, score entailment (003333): 00 = Not entailed/Contradicted, $1$ = Related not entailed, $2$ = Mostly entailed, $3$ = Perfectly entailed.
@@
[2] Chain-of-Thought Decomposed Reasoning
@@
PHASE 2 – Fine-grained Criterion Scoring:
@@
System prompt when Phase 1 = "Yes": You are a rater evaluating entailment. Provide a final score ($2$ or $3$): $2$ = Mostly entailed, $3$ = Perfectly entailed. Your output MUST be JSON: {"relevance_score": <2_or_3>}
System prompt when Phase 1 = "No": You are a rater evaluating entailment. Provide a final score (00 or $1$): 00 = Not entailed/Contradicted, $1$ = Related but not entailed. Your output MUST be JSON: {"relevance_score": <0_or_1>}
@@
User prompt when Phase 1 = "No": Hypothesis is not entailed. Rate how (00 or $1$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
[3] Multi-Dimensional Attribute Scoring
PHASE 1 – Sub-criterion Scoring:
System prompt: Please assess how well the hypothesis is entailed by the premise according to the following criteria. Use a score (003333): 00 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion. $2$ = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
@@
System prompt: You are a search quality rater evaluating entailment. Given a premise, hypothesis, and sub-scores, provide a final score (003333): $3$ = Perfectly entailed, $2$ = Mostly entailed, $1$ = Related but not entailed, 00 = Not entailed/Contradicted. Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
@@
[1] Zero-shot Direct Scoring for STS
System prompt: You are a semantic-similarity rater. Output JSON with a score (005555) where 00 = unrelated, $5$ = semantically equivalent.
User prompt: Rate the semantic similarity 005555: 00 = unrelated $\mid$ $1$ = slight $\mid$ $2$ = partial $\mid$ $3$ = moderate $\mid$ $4$ = high $\mid$ $5$ = equivalent
@@
granite-embedding-125m-english 125M Medium RoBERTa 768 14 encoder
bge-large-en-v1.5 335M Large 1024 15 encoder
mxbai-embed-large-v1 335M Large 1024 16 encoder
snowflake-arctic-embed-s 33M Small e5-unsupervised-small 384 17 encoder
bge-base-en-v1.5 109M Medium 768 18 encoder
bge-small-en-v1.5 33M Small 384 19 encoder
multilingual-e5-large-instruct 560M Large xlm-roberta-large 1024 20 encoder
thenlper-gte-base 109M Medium EBRT-base 768 21 encoder
multilingual-e5-large 560M Large xlm-roberta-large 1024 22 encoder
gte-Qwen2-1.5B-instruct 1.5B XL Qwen2-1.5B 1536 23 decoder
e5-base-v2 109M Medium bert-large-uncased 1024 24 encoder
e5-large-v2 335M Large bert-base-uncased 768 25 encoder
e5-small-v2 33M Small MiniLM 384 26 encoder
all-mpnet-base-v2 109M Medium microsoft/mpnet-base 768 27 encoder
gte-base-en-v1.5 137M Medium EBRT-base 768 28 encoder
@@
NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2 cc-by-nc-4.0
@@
Linq-Embed-Mistral https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral cc-by-nc-4.0
NV-Embed-v1 https://huggingface.co/nvidia/NV-Embed-v1 cc-by-nc-4.0
SFR-Embedding-Mistral https://huggingface.co/Salesforce/SFR-Embedding-Mistral cc-by-nc-4.0
@@
all-mpnet-base-v2 https://huggingface.co/sentence-transformers/all-mpnet-base-v2 apache-2.0
@@
llmrails-ember-v1 https://huggingface.co/llmrails/ember-v1 mit
@@
Biological 12.93 19.75 19.16 25.18 18.82 34.19 21.67
Chemical 14.74 20.20 19.57 25.92 15.82 26.73 20.50
Environmental 14.03 26.03 23.66 21.48 19.75 34.39 23.23
Extraterrestrial 13.13 21.89 15.62 27.12 17.38 30.09 20.87
Geohazard 14.68 17.35 18.51 21.79 15.51 34.44 20.38
Meteorological&hydrological 12.63 19.33 23.63 22.10 19.47 30.22 21.23
Societal 14.51 29.00 17.56 24.04 18.59 22.57 21.05
Technological 15.20 24.16 19.15 23.58 18.63 37.00 22.95
Avg. 13.98 22.21 19.61 23.90 18.00 31.20 21.48
@@
Biological 62.58 37.35 43.67 27.64 45.76 71.38 48.06
Chemical 64.42 38.19 41.06 31.31 46.84 70.43 48.71
Environmental 62.58 38.36 41.83 29.77 48.23 71.06 48.64
Extraterrestrial 58.55 34.42 37.55 31.45 46.54 73.40 46.99
Geohazard 61.12 36.31 40.13 28.31 47.12 68.06 46.84
Meteorological&hydrological 62.22 39.15 42.08 26.14 45.62 70.57 47.63
Societal 66.77 47.63 43.10 31.95 51.06 72.05 52.09
Technological 61.51 41.51 41.80 33.27 47.08 70.86 49.34
Avg. 62.47 39.12 41.40 29.98 47.28 70.98 48.54
@@
Biological 60.86 54.84 46.25 53.77 42.53 70.95 54.87
Chemical 60.99 55.23 43.96 53.63 37.30 70.23 53.56
Environmental 62.02 55.77 47.11 51.32 43.30 71.72 55.21
Extraterrestrial 59.48 53.38 49.06 53.30 37.65 72.35 54.20
Geohazard 60.10 54.26 45.96 49.82 36.83 70.45 52.90
Meteorological&hydrological 59.98 58.31 47.26 51.44 39.13 70.93 54.51
Societal 61.99 59.05 48.04 53.11 42.88 71.83 56.15
Technological 59.83 57.66 48.85 53.55 39.39 71.68 55.16
Avg. 60.66 56.06 47.06 52.49 39.88 71.27 54.57

### v2

@@ Acknowledgments
This work used ACES at TAMU, DeltaAI and Delta GPU at the National Center for Supercomputing Applications through allocation CIV250019 and CIV250021 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Structural PDF File Processing Pipeline
Model Bio Chem Env Extra Geo MH Soc Tech
Linq-Embed-Mistral 68.07 66.75 67.82 67.30 65.40 67.13 68.56 67.36
SFR-Embedding-Mistral 67.74 66.70 68.03 66.43 65.40 67.03 68.15 67.37
inf-retriever-v1 67.75 66.40 66.92 65.04 65.52 66.66 68.15 66.75
inf-retriever-v1-1.5b 65.71 64.45 65.36 63.33 64.31 65.45 66.59 65.54
NV-Embed-v2 65.23 64.51 65.55 64.79 63.51 64.94 65.95 64.89
gte-Qwen2-1.5B-instruct 65.15 64.35 64.94 62.56 63.16 63.97 66.12 65.16
multilingual-e5-large 63.00 62.91 64.01 62.11 62.59 63.10 64.26 64.10
e5-mistral-7b-instruct 63.33 62.13 63.49 62.25 60.88 62.70 63.85 63.04
multilingual-e5-large-instruct 62.64 61.79 62.76 62.48 60.49 62.70 63.60 62.95
e5-small-v2 62.73 62.56 64.17 60.59 60.71 62.68 62.29 62.29
e5-base-v2 61.86 61.34 63.05 60.97 61.08 61.96 62.98 61.75
e5-large-v2 60.09 59.49 60.84 59.77 59.18 60.48 61.15 60.74
NV-Embed-v1 61.37 58.86 61.91 58.84 59.13 61.13 61.23 60.66
granite-embedding-125m 58.00 57.52 58.09 57.21 56.80 57.89 58.32 57.61
arctic-embed-m-v2.0 56.47 55.78 56.14 55.05 54.88 55.66 55.98 55.29
mxbai-embed-large-v1 56.18 55.38 55.96 54.77 54.56 55.47 55.88 55.07
gte-base-en-v1.5 54.78 54.29 55.49 53.19 52.99 54.19 54.60 53.98
bge-base-en-v1.5 53.57 52.78 54.18 52.38 51.98 53.18 53.77 52.88
gte-large-en-v1.5 63.23 62.54 63.86 62.32 61.92 63.05 64.10 63.07
snowflake-arctic-embed-l-v2.0 53.86 53.09 54.44 52.78 52.36 53.58 53.96 53.17
bge-large-en-v1.5 53.18 52.49 53.88 52.08 51.78 52.79 53.49 52.58
gte-Qwen2-7B-instruct 52.90 52.20 53.70 51.90 51.60 52.60 53.20 52.30
bge-small-en-v1.5 45.05 44.45 35.75 44.91 44.14 45.23 45.95 44.73
snowflake-arctic-embed-s 40.91 40.51 43.60 38.26 38.71 40.15 41.65 40.50
snowflake-arctic-embed-m-v1.5 42.30 42.79 44.14 40.49 41.20 41.99 42.47 42.19
snowflake-arctic-embed-l 35.15 35.05 37.77 34.28 32.53 33.81 35.92 34.59
thenlper-gte-base 35.84 35.28 33.35 32.19 31.19 32.71 34.83 33.84
snowflake-arctic-embed-m 30.72 31.00 32.51 31.40 29.45 30.71 31.62 30.74
snowflake-arctic-embed-m-long 30.82 30.67 32.16 30.77 29.44 30.00 31.87 30.75
thenlper-gte-small 31.92 31.53 28.83 30.25 27.72 30.26 29.96 31.27
Table 4: Performance of 30 open-source retrieval models under the exact-search setting across different disaster event types. Each cell shows the mean NDCG@10 over the 6 search tasks for that event type.
@@
All disaster management-related data (in PDF format) is obtained from publicly available sources with no personally identifiable information. Hence, explicit consent was not required. We chose PDFs because they typically contain more structured, information-rich, and credible content, often originating from peer-reviewed publications or official institutions. PDF files are collected using googlesearch-python (v1.3.0) and processed with PyMuPDF (v1.24.10) for content extraction. The extracted PDFs are then processed into text chunks through the following steps:
@@
Since our benchmark emphasizes retrieval performance across event types and search intents rather than source analysis, we summarize the corpus by the number of PDFs per general event type. Each general event type includes multiple specific event types, which were used as search queries. The number of collected PDFs is thus correlated with the number of associated specific types, as shown in Table 5.
General event type Number of PDFs (percentage)
MH 1683 (19.89%)
Bio 2343 (27.68%)
Geo 868 (10.26%)
Tech 1598 (18.88%)
Extra 254 (3.00%)
Chem 767 (9.06%)
Soc 234 (2.76%)
Env 717 (8.47%)
Table 5: Distribution of event types by number of PDFs
Appendix B Evaluation of LLM-generated passages
B.1 Evaluation of LLM document summaries
Experiment design.
We evaluate whether each LLM-generated summary accurately captured the main content of the original material, which motivated their inclusion in the corpus. A total of 48 summaries (six per general event type) are randomly sampled. Three PhD students independently annotate each summary for fluency and content accuracy, following detailed guidelines adapted from Song et al. (2024).
Results and analysis.
Fluency is defined as the percentage of summaries judged grammatically and semantically well-formed, while content accuracy measures whether summaries captured the main ideas of the source material. The average results show 100% fluency and 96.88% content accuracy. The Fleiss’ Kappa score of 0.793 indicated substantial inter-annotator agreement, supporting the reliability of the evaluation. Overall, LLM-generated summaries are fluent and highly faithful to the original documents.
B.2 Evaluation of LLM-generated passages from page content
Experiment design.
We further assess LLM-generated passages created directly from page content, focusing on their fluency and adherence to style requirements. The QAdoc intent is excluded, as its format (document summaries) had already been validated. We randomly sampled 40 passages (eight per intent across five intents) and asked three PhD students to evaluate them. The evaluators followed the same style guidelines provided to the LLM during generation.
Results and analysis.
LLM-generated passages achieved an average fluency score of 98.75% and a style compliance score of 92.5%. The Fleiss’ Kappa of 0.760 again confirmed substantial agreement among annotators. These results indicate that the generated passages are both fluent and well aligned with intent-specific stylistic requirements.
B.3 Role of noisy passages in realistic IR evaluation
While our evaluations confirm the overall quality of LLM-generated content, we emphasize the value of including imperfect or noisy passages in the corpus. Real-world IR corpora naturally contain irrelevant or off-topic data, and such noise can enhance the realism of evaluations. Introducing noisy passages allows us to test the robustness of IR models by assessing whether they can correctly identify irrelevant content and assign low relevance scores. For example, a generated tweet unrelated to a query serves as a negative case; an effective IR model should detect this mismatch and rank it accordingly.
Appendix C Prompt Templates for Query Generation
Prompts for query generation based on disaster management-related passage under different search intents for QA, QAdoc, Twitter, FC, NLI, STS are in Tables 7, 8, 9, 10, 11, and 12.
Appendix D Prompt Templates for Relevance Labeling
This section presents the prompt templates used for LLM-based relevance judgments across six search intents, employing three prompting strategies: Zero-shot Direct Scoring, Chain-of-Thought Decomposed Reasoning, and Multi-Dimensional Attribute Scoring. For QA, QAdoc, and Twitter tasks, we adapt templates from Thomas et al. (2024); Farzi and Dietz (2024), as shown in Table 13. Based on these templates, we design relevance prompts for FC, NLI, and STS tasks, shown in Tables 14, 15, and 16, respectively. For STS, we adopt only Zero-shot Direct Scoring, as our preliminary experiments show it yields higher agreement with human labels (Cohen’s kappa). The estimated cost of generating 9,600 user queries and labeling over 1.3 million query-passage using GPT-4o-mini API is about \$1,400.
Appendix E Additional Analyses of Labeled query-passage pairs
@@
Appendix F LVHL Dataset Construction
@@
Appendix G LVHQ Dataset Construction
We sample 48 domain passages developed in Section 3.3, ensuring one passage per retrieval task and keeping all sampled passages different from those used in developing DisastIR. For each passage, a domain expert in the disaster management field is asked to read the passage and write a realistic user query that reflects a practical information need based on the content, resulting in 48 human-authored queries. The Human expert is given the same instructions for the query written (shown in Tables 7, 8, 9, 10, 11, 12) as those given to LLM to ensure fair comparison. In parallel, for the same set of passages, we also generate 48 queries using LLM in the same way as described in Section 3.4.
@@ Appendix G LVHQ Dataset Construction
Appendix H Information of Evaluated Models and Model Implementation
@@
Appendix I Performance of Evaluated Models
Performance of all evaluated models in all 48 search tasks in DisastIR is shown in Tables 22, 23, 24, 25, and 26.
Appendix J Additional Analyses of Model Performance across 48 Tasks
NV-Embed-v2 achieves the best performance on all NLI-related tasks (See Table 3 and Figure 4 in the main content). However, as shown in Figure 7, its poor results on Twitter-related tasks significantly lower its overall performance in DisastIR. This reflects its limitation in handling informal, noisy, and contextually ambiguous nature of social media content. Given the importance of Twitter as a real-time, crowd-sourced information source during disasters (Alam et al., 2021; Yin et al., 2024; Lei et al., 2025), this weakness raises concerns about its reliability in real-world disaster response scenarios.
All four models perform poorly on NLI-related tasks, with the best achieving only an average score of 58.39 (Figure 7). Further analysis of outliers in the box plot (See Figure 3 in the main content) reveals that tasks causing significant performance drops consistently involve NLI search intents (Figure 8). This reveals a key limitation of current open-source SOTA retrievers, that they struggle with the complex reasoning required for NLI tasks in disaster contexts. Such limitations may lead to incorrect results or failure to retrieve critical information, which can negatively impact decision-making in disaster situations.
Reference Labeling Cohen’s $\kappa$
Annotator 1 0.695
Annotator 2 0.681
Annotator 3 0.734
Rounded average 0.690
Majority vote (min in tie) 0.803
Majority vote (max in tie) 0.715
Fleiss’ $\kappa$ (inter-annotator) 0.777
Table 6: Agreement between LLM-based relevance labels and human annotations on the in-domain sample (96 query–passage pairs).
@@ Appendix I Performance of Evaluated Models
System prompt: You are a search quality rater evaluating overall relevance. Given a query, passage, and sub‐scores, provide a final score (0–3): 3 = Perfectly relevant, 2 = Highly relevant, 1 = Related, 0 = Irrelevant. Your output MUST be JSON: {"final_relevance_score": <0-3>}
User prompt: Please rate how relevant the passage is based on the given sub‐scores.Query: {query}, Passage: {passage}, Exactness: LLM score, Coverage: LLM score, Topicality: LLM score, Contextual Fit: LLM score.
@@
[1] Zero‐shot Direct Scoring
@@
User prompt: Given a claim and a passage, score relevance (033): 0 = Irrelevant, $1$ = Related (no help), $2$ = Relevant (unclear/mixed), $3$ = Direct support/refutation.
@@
[2] Chain‐of‐Thought Decomposed Reasoning
@@
PHASE 2 – Fine‐grained Criterion Scoring:
System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim. Use the following scoring scale (033): 0 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion.22 = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
@@
System prompt when Phase 1 = "Yes": You are a rater evaluating evidence for fact-checking. Score ($2$ or $3$): 22 = Highly relevant (some support/refutation), $3$ = Perfectly relevant (direct support/refutation). Your output MUST be JSON: {"relevance_score": <2_or_3>} System prompt when Phase 1 = "No": You are a rater evaluating evidence for fact-checking. Score (0 or $1$): 0 = Irrelevant, $1$ = Related (no support/refutation). Your output MUST be JSON: {"relevance_score": <0_or_1>}
@@
Passage irrelevant for evidence. Rate its relevance (0 or $1$). Claim: {Query}, Passage: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
[3] Multi‐Dimensional Attribute Scoring
PHASE 1 – Sub‐criterion Scoring:
System prompt: Please assess how well the provided passage serves as evidence for evaluating the claim according to the following criteria. Use a score (033): 0 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion. $2$ = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
@@
System prompt: You are a search quality rater evaluating evidence relevance for fact-checking. Given a claim, passage, and sub‐scores, provide a final score (033): 33 = Perfectly relevant (direct support/refutation), $2$ = Highly relevant (helps verification), $1$ = Related (topic match, no verification aid), 0 = Irrelevant. Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
@@
[1] Zero‐shot Direct Scoring
@@
User prompt: Given a premise and hypothesis, score entailment (033): 0 = Not entailed/Contradicted, $1$ = Related not entailed, $2$ = Mostly entailed, $3$ = Perfectly entailed.
@@
[2] Chain‐of‐Thought Decomposed Reasoning
@@
PHASE 2 – Fine‐grained Criterion Scoring:
@@
System prompt when Phase 1 = "Yes": You are a rater evaluating entailment. Provide a final score ($2$ or $3$): 22 = Mostly entailed, $3$ = Perfectly entailed. Your output MUST be JSON: {"relevance_score": <2_or_3>}
System prompt when Phase 1 = "No": You are a rater evaluating entailment. Provide a final score (0 or $1$): 0 = Not entailed/Contradicted, $1$ = Related but not entailed. Your output MUST be JSON: {"relevance_score": <0_or_1>}
@@
User prompt when Phase 1 = "No": Hypothesis is not entailed. Rate how (0 or $1$). Premise: {Query}, Hypothesis: {Passage}. Output JSON: {"relevance_score": <0_or_1>}
[3] Multi‐Dimensional Attribute Scoring
PHASE 1 – Sub‐criterion Scoring:
System prompt: Please assess how well the hypothesis is entailed by the premise according to the following criteria. Use a score (033): 0 = Not relevant at all / No information provided. $1$ = Marginally relevant / Partially addresses the criterion. $2$ = Fairly relevant / Adequately addresses the criterion. $3$ = Highly relevant / Fully satisfies the criterion. Your output MUST be JSON: {"criterion_score": <0_to_3>}
@@
System prompt: You are a search quality rater evaluating entailment. Given a premise, hypothesis, and sub‐scores, provide a final score (033): 33 = Perfectly entailed, $2$ = Mostly entailed, $1$ = Related but not entailed, 0 = Not entailed/Contradicted. Your output MUST be JSON: {"final_relevance_score": <0_to_3>}
@@
[1] Zero‐shot Direct Scoring for STS
System prompt: You are a semantic‐similarity rater. Output JSON with a score (055) where 0 = unrelated, $5$ = semantically equivalent.
User prompt: Rate the semantic similarity 055: 0 = unrelated $\mid$ 11 = slight $\mid$ 22 = partial $\mid$ 33 = moderate $\mid$ 44 = high $\mid$ 55 = equivalent
@@
snowflake-arctic-embed-m-long 137M Medium nomic-embed-text-v1-unsupervised 768 14 encoder
granite-embedding-125m-english 125M Medium RoBERTa 768 15 encoder
bge-large-en-v1.5 335M Large 1024 16 encoder
mxbai-embed-large-v1 335M Large 1024 17 encoder
snowflake-arctic-embed-s 33M Small e5-unsupervised-small 384 18 encoder
bge-base-en-v1.5 109M Medium 768 19 encoder
bge-small-en-v1.5 33M Small 384 20 encoder
multilingual-e5-large-instruct 560M Large xlm-roberta-large 1024 21 encoder
thenlper-gte-base 109M Medium EBRT-base 768 22 encoder
multilingual-e5-large 560M Large xlm-roberta-large 1024 23 encoder
thenlper-gte-small 33M Small MiniLM-L12-H384 384 24 encoder
gte-Qwen2-1.5B-instruct 1.5B XL Qwen2-1.5B 1536 25 decoder
e5-base-v2 109M Medium bert-large-uncased 1024 26 encoder
e5-large-v2 335M Large bert-base-uncased 768 27 encoder
e5-small-v2 33M Small MiniLM 384 28 encoder
gte-base-en-v1.5 137M Medium EBRT-base 768 29 encoder
@@
NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2 $c$-by-nc-4.0
@@
Linq-Embed-Mistral https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral $c$-by-nc-4.0
NV-Embed-v1 https://huggingface.co/nvidia/NV-Embed-v1 $c$-by-nc-4.0
SFR-Embedding-Mistral https://huggingface.co/Salesforce/SFR-Embedding-Mistral $c$-by-nc-4.0
@@
snowflake-arctic-embed-m https://huggingface.co/Snowflake/snowflake-arctic-embed-m-long apache-2.0
@@
thenlper-gte-small https://huggingface.co/thenlper/gte-small mit
@@
Biological 70.58 60.91 68.01 61.43 55.36 74.59 65.15
Chemical 70.02 58.58 64.21 64.40 55.48 73.43 64.35
Environmental 69.88 56.90 65.41 64.73 59.01 73.74 64.94
Extraterrestrial 67.56 56.18 61.33 62.44 53.14 74.72 62.56
Geohazard 70.30 58.34 63.81 59.60 54.98 71.91 63.16
Meteorological&hydrological 70.45 59.65 66.28 59.56 54.95 72.93 63.97
Societal 71.41 61.67 67.07 64.68 58.40 73.51 66.12
Technological 69.44 61.46 65.57 65.90 54.54 74.01 65.16
Avg. 69.96 59.21 65.21 62.84 55.73 73.61 64.43
@@
Biological 70.42 47.68 64.56 32.33 53.79 75.93 57.45
Chemical 71.77 47.98 62.94 35.13 52.81 75.01 57.61
Environmental 69.36 44.00 63.04 30.11 56.15 75.33 56.33
Extraterrestrial 67.12 44.91 59.87 32.20 50.51 74.59 54.87
Geohazard 71.17 47.91 63.11 30.85 53.78 74.34 56.86
Meteorological&hydrological 70.30 48.08 64.94 27.68 52.61 73.10 56.12
Societal 71.81 50.68 64.35 32.02 57.89 74.82 58.59
Technological 70.45 49.94 63.06 34.64 53.53 75.80 57.90
Avg. 70.30 47.65 63.24 31.87 53.88 74.86 56.97
@@
Biological 60.71 54.57 45.27 53.77 42.13 70.02 54.41
Chemical 60.83 54.93 43.47 53.53 37.44 69.37 53.26
Environmental 61.88 55.46 46.43 51.08 43.30 71.11 54.88
Extraterrestrial 59.20 53.20 48.45 53.15 37.86 71.54 53.90
Geohazard 60.00 54.10 45.55 49.61 37.01 69.57 52.64
Meteorological&hydrological 59.74 58.13 46.70 51.29 39.06 69.79 54.12
Societal 61.78 58.93 47.54 52.86 42.46 70.98 55.76
Technological 59.54 57.48 48.11 53.45 39.55 70.89 54.84
Avg. 60.46 55.85 46.44 52.34 39.85 70.41 54.23